I am only stuck on part C. I have tried two answers that I got and they aren't correct.A baseball player throws a baseball when it is 5 ft above the ground. The ball leaves his hand at an angle of 16° with respec
Answer #1

Gfven 16 fefSec Jnitial Velocity Vo156 1 ft = 0-305 metre that Krous - 47.5488 m metse 156ft=156 x 0-305 the =1-524 metre 5X0